The mixed fraction  5 whole 4 by 7 can be expressed as

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


5 whole 4 by 7can be expressed as = 5 + (4/7)

= (35 + 4)/7
= 39/7

Thanks